# Nonlinear morphoelastic energy based theory for stimuli responsive elastic shells

### v2

@@ Appendix B Full form of the reduced energies
In this sections we present the full form of the reduced energy, when the through-the-thickness non-elastic stimulus in includes also $\bar{\rho}_{2}\neq 0$ and $\bar{\rho}_{3}\neq 0$, other than the contribution of $\bar{\rho}_{1}$.
@@ B.1 Neo-Hookean material
$\Psi_{nc,2d}=\int_{\mathcal{S}_{0}}\left[h_{0}\bar{\rho}_{1}\left(\frac{\mu}{2}w^{\textup{NH}}_{s}+p(J_{s}-1)\right)+\frac{h_{0}^{3}}{12}\frac{1}{2\bar{\rho}_{1}^{3}}\left(\mu w^{\textup{NH}}_{b}+2\bar{\rho}_{1}^{3}pf_{p}\right)\right]\sqrt{\det(\bar{\mathrm{a}}_{\alpha\beta})}\,\mathrm{d}\eta_{1}\mathrm{d}\eta_{2}+O(h^{5})\,$, (94)
where the integral is evaluated with respect to the coordinates in the reference configuration $\mathcal{S}_{0}$ and $h_{0}$ indicates the undeformed thickness. Symbol $\mu$ indicates the shear modulus and $p$ is the Lagrange multiplier that enforces incompressibility; $J_{s}$ is the quantity that indicates the area change of the mid-surface and the dimensional reduction is performed assuming that $J_{s}=1$. The stretching and bending terms are given by:
$w^{\textup{NH}}_{s}$ $=\mathrm{a}_{\alpha\beta}\bar{\mathrm{a}}^{\alpha\beta}-2,$ (95)
$w^{\textup{NH}}_{b}$ $=\bar{\rho}_{1}^{3}f_{1}w_{s}^{NH}+2\bar{\rho}_{1}^{4}f_{2}\mathrm{b}_{\alpha\beta}\bar{\mathrm{a}}^{\alpha\beta}+2\bar{\rho}_{1}^{4}f_{3}\mathrm{a}_{\alpha\beta}\bar{\mathrm{b}}^{\alpha\beta}$
$+\bar{\rho}_{1}^{6}\left(-4\mathrm{b}_{\alpha\beta}\bar{\mathrm{b}}^{\alpha\beta}+\mathrm{c}_{\alpha\beta}\bar{\mathrm{a}}^{\alpha\beta}+3\mathrm{a}_{\alpha\beta}\bar{\mathrm{c}}^{\alpha\beta}\right)+\bar{\rho}_{1}^{2}f_{4}\,$, (96)
@@
$f_{1}$ $=\bar{\rho}_{1}^{3}\bar{K}-6\bar{\rho}_{1}\bar{\rho}_{2}\bar{H}+3\bar{\rho}_{3},$ (97)
$f_{2}$ $=2\bar{\rho}_{1}^{2}\bar{H}-\bar{\rho}_{1}+\bar{\rho}_{2}\left(\bar{\rho}_{2}(H-\bar{H})-2\right),$ (98)
$f_{3}$ $=3\bar{\rho}_{2}-2\bar{\rho}_{1}^{2}\bar{H},$ (99)
$f_{4}$ $=2\bar{\rho}_{1}^{4}\left(6H(H-\bar{H})-K+\bar{K}\right)-8\bar{\rho}_{1}^{3}\bar{H}+4\bar{\rho}_{2}^{2}\left(\bar{\rho}_{2}(H-\bar{H})+1\right)^{2}$ (100)
$-8\bar{\rho}_{1}\bar{\rho}_{2}\left(\bar{\rho}_{2}(H-\bar{H})+1\right)+4\bar{\rho}_{1}^{2}\left(\bar{\rho}_{2}H(2\bar{\rho}_{2}\bar{H}-3)+\bar{\rho}_{2}\bar{H}(5-2\bar{\rho}_{2}\bar{H})+1\right),$
$f_{p}$ $=\bar{\rho}_{1}^{3}(J_{s}-1)\bar{K}$
$+6\bar{\rho}_{1}\left(J_{s}(H-\bar{H})\left(H\left(\bar{\rho}_{1}^{2}+\bar{\rho}_{2}^{2}\right)-\bar{\rho}_{2}\right)-\bar{\rho}_{1}HJ_{s}+\bar{\rho}_{2}\bar{H}\right)+3\bar{\rho}_{3}(J_{s}-1)\,.$ (101)
@@ B.2 Ciarlet-Geymonat material
$\left(\mu^{*}\right)^{-1}\Psi_{cg}=\int_{\mathcal{S}_{0}}\left[h_{0}w^{CG}_{s}+\frac{h_{0}^{3}}{12}w^{CG}_{b}\right]\sqrt{\det(\bar{\mathrm{a}}_{\alpha\beta})}\,\mathrm{d}\eta_{1}\mathrm{d}\eta_{2}+O(h^{5})\,$, (102)
where we define the rescaled bulk modulus as $\lambda=\lambda^{*}/\mu^{*}$. The stretching contribution $w^{\textup{CG}}_{s}$ and the function $\rho_{1}$ are already presented in the main text as they do not depend on the higher order terms in the through-the-thickness $\bar{\rho}_{2}$ and $\bar{\rho}_{3}$ reads
The bending contribution, including the coefficients $\bar{\rho}_{2}$ and $\bar{\rho}_{3}$ is given by
@@
$g_{1}$ $=\bar{\rho}_{1}^{3}\bar{K}-6\bar{\rho}_{1}\bar{\rho}_{2}\bar{H}+3\bar{\rho}_{3}$ (104)
$g_{2}$ $=\left(2\rho_{1}\bar{\rho}_{1}^{2}\bar{H}-2\rho_{1}\bar{\rho}_{2}-\bar{\rho}_{1}\rho_{2}\right)$ (105)
$g_{3}$ $=\bar{\rho}_{1}\left(3\bar{\rho}_{2}-2\bar{\rho}_{1}^{2}\bar{H}\right)$ (106)
@@
Minimization of (103) with respect to $\rho_{2}$ leads to:
$\rho_{2}=\frac{1}{2}\bar{\rho}_{1}^{2}\left[\left(\frac{H(\lambda+2)(2J_{s}\lambda-1)}{(J_{s}\lambda+2)^{2}}-\frac{2\bar{H}J_{s}\lambda\sqrt{\lambda+2}}{(J_{s}\lambda+2)^{3/2}}\right)+\frac{\mathrm{b}_{\alpha\beta}\bar{\mathrm{a}}^{\alpha\beta}}{4J_{s}\lambda+8}\right]+\bar{\rho}_{2}\frac{\sqrt{\lambda+2}}{\sqrt{J_{s}\lambda+2}}.$ (108)

### v3

@@ Appendix B Full form of the reduced energies
In this sections we present the full form of the reduced energies, when the through-the-thickness non-elastic stimulus includes also $\bar{\rho}_{2}\neq 0$ and $\bar{\rho}_{3}\neq 0$, other than the contribution of $\bar{\rho}_{1}$.
@@ B.1 Neo-Hookean material
$\Psi_{nc,2d}=\int_{\mathcal{S}_{0}}\left[h_{0}w^{\textup{NH}}_{s}+\frac{h_{0}^{3}}{12}w^{\textup{NH}}_{b}\right]\sqrt{\det(\bar{\mathrm{a}}_{\alpha\beta})}\,\mathrm{d}\eta^{1}\mathrm{d}\eta^{2}+O(h^{5})\,$, (99)
where the integral is evaluated with respect to the coordinates in the reference configuration $\mathcal{S}_{0}$ and $h_{0}$ indicates the undeformed thickness. The stretching contribution $w^{\textup{NH}}_{s}$ depends only on $\bar{\rho}_{1}$ so it remains equal to what has been presented in the main text, while the bending term in the iso-areal case is modified by $\bar{\rho}_{1}$ and $\bar{\rho}_{2}$ as
$w^{\textup{NH}}_{b}=$ $\frac{\mu^{*}}{2}\left(2\frac{\bar{K}}{\bar{\rho}_{1}}\left(\bar{\rho}_{1}^{3}\bar{K}-6\bar{\rho}_{1}\bar{\rho}_{2}\bar{H}+3\bar{\rho}_{3}\right)w^{\textup{NH}}_{s}+2\bar{\rho}_{1}\left(\bar{\rho}_{1}^{2}(H+\bar{H})-3\bar{\rho}_{2}\right)\left(H-\bar{H}\right)\mathrm{b}_{\alpha\beta}\bar{\mathrm{a}}^{\alpha\beta}\right$.
$\left.+2\bar{\rho}_{1}\left(3\bar{\rho}_{2}-2\bar{\rho}_{1}^{2}\bar{H}\right)\bar{H}\mathrm{a}_{\alpha\beta}\bar{\mathrm{b}}^{\alpha\beta}+p_{0}f_{p0}+p_{1}f_{p1}\right$. (100)
$\left.+\bar{\rho}_{1}^{3}\left(-4\mathrm{b}_{\alpha\beta}\bar{\mathrm{b}}^{\alpha\beta}+\mathrm{c}_{\alpha\beta}\bar{\mathrm{a}}^{\alpha\beta}+3\mathrm{a}_{\alpha\beta}\bar{\mathrm{c}}^{\alpha\beta}\right)+f_{1}\right)\,$, (101)
@@
$f_{1}$ $=2\bar{\rho}_{1}\left(\bar{\rho}_{1}^{2}(2(H-\bar{H})(4H+\bar{H})-K+\bar{K})+6\bar{\rho}_{2}(\bar{H}-H)\right),$ (102)
$f_{p0}$ $=\frac{24}{\mu^{*}}\bar{\rho}_{1}J_{s}(H-\bar{H})\left(\bar{\rho}_{1}^{2}H-\bar{\rho}_{2}\right)\,$ (103)
$f_{p1}$ $=\frac{4}{\mu^{*}}\bar{\rho}_{1}^{2}(-2HJ_{s}+\bar{H}J_{s}+\bar{H})+\bar{\rho}_{2}(J_{s}-1)\,.$ (104)
@@ B.2 Ciarlet-Geymonat material
$\left(\mu^{*}\right)^{-1}\Psi_{cg}=\int_{\mathcal{S}_{0}}\left[h_{0}w^{CG}_{s}+\frac{h_{0}^{3}}{12}w^{CG}_{b}\right]\sqrt{\det(\bar{\mathrm{a}}_{\alpha\beta})}\,\mathrm{d}\eta^{1}\mathrm{d}\eta^{2}+O(h^{5})\,$, (105)
where we define the rescaled bulk modulus as $\lambda=\lambda^{*}/\mu^{*}$. The stretching contribution $w^{\textup{CG}}_{s}$ has been already presented in the main text because it does not depend on the higher order terms in the through-the-thickness, $\bar{\rho}_{2}$ and $\bar{\rho}_{3}$. The bending contribution is instead modified as
@@
$g_{1}$ $=\bar{\rho}_{1}^{3}\bar{K}-6\bar{\rho}_{1}\bar{\rho}_{2}\bar{H}+3\bar{\rho}_{3}\,$, (107)
$g_{2}$ $=\left(2\rho_{1}\bar{\rho}_{1}^{2}\bar{H}-2\rho_{1}\bar{\rho}_{2}-\bar{\rho}_{1}\rho_{2}\right)\,$, (108)
$g_{3}$ $=\bar{\rho}_{1}\left(3\bar{\rho}_{2}-2\bar{\rho}_{1}^{2}\bar{H}\right)\,$, (109)
@@
Minimization of eq. (106) with respect to $\rho_{2}$ leads to:
$\rho_{2}=\frac{1}{2}\bar{\rho}_{1}^{2}\left[\left(\frac{H(\lambda+2)(2J_{s}\lambda-1)}{(J_{s}\lambda+2)^{2}}-\frac{2\bar{H}J_{s}\lambda\sqrt{\lambda+2}}{(J_{s}\lambda+2)^{3/2}}\right)+\frac{\mathrm{b}_{\alpha\beta}\bar{\mathrm{a}}^{\alpha\beta}}{4J_{s}\lambda+8}\right]+\bar{\rho}_{2}\frac{\sqrt{\lambda+2}}{\sqrt{J_{s}\lambda+2}}\,$, (111)
with $\rho_{1}$ as presented in the main text.